Find the slope and y-intercept for each line that has a slope
and y-intercept.
x + 2y = 3

Answer by angelagallegos(8)

The first thing you want to do is to solve for y. putting it in slope intercept form y=mx+b
x+2y= 3 subtract x from both sides
-x -x x will cancel
2y=-x+3 divide both sides by 2
y= -x+3
2 make the fractions
y=-1/2x+3/2 any number with the x will be the slope and the b is y-intercept
answer= slope= -1/2 and Y-intercept= 3/2